Differentiate the following functions.(use Chain Rules)




1. y=(5x² —2x + 1)—³

1
Expert's answer
2021-12-05T18:20:27-0500

Let us differentiate the function


1. "y=(5x^2 -2x + 1)^{-3}"


It follows that


"y'=-3(5x^2 -2x + 1)^{-4}(5x^2 -2x + 1)'=-3(5x^2 -2x + 1)^{-4}(10x -2)"


"=-6(5x -1)(5x^2 -2x + 1)^{-4}."